# Structure-Dynamics Correlation and Its Link to Fragility and Dynamic Heterogeneity

### v2

@@ IV.1 The temperature and SOP dependence of the energy barrier and its correlation with fragility
and
$1/\Delta E^{ma}=\frac{K_{VFT}}{T_{VFT}}-\gamma.K_{VFT}.\frac{1}{\Delta((\beta\Phi)^{-1})}=A^{\prime}+B^{\prime}.\frac{1}{\Delta((\beta\Phi)^{-1})}$ (22)
From Eq.21 we get A = $\frac{T_{VFT}}{K_{VFT}}$ and B = $\gamma.\frac{T_{VFT}^{2}}{K_{VFT}}$ and the slope is again not proportional to but inversely proportional to the fragility. Thus similar to the argument given for the temperature dependence of the energy barrier, we find that as shown in Fig.1(left panels) and Fig.2 (left panels) the slopes do not always follow fragility. In cases where it follows fragility, it is because for more fragile systems the $T_{VFT}$ is higher. From Eq.22 we get that $A^{\prime}=\frac{K_{VFT}}{T_{VFT}}$ and $B^{\prime}=\gamma.K_{VFT}$. Thus, similar to that observed in temperature dependence, the slope of $\frac{1}{\Delta E^{ma}}$ is proportional to fragility. However, in this case, there is this additional factor $\gamma$ which describes the dependence of the SOP on the temperature.
@@
Note that the temperature dependence of the energy barrier, $\Delta E^{ma}$ comes from the temperature dependence of the SOP, and the energy barrier has a temperature independent term. A similar observation can also be made for $\frac{1}{\Delta E^{ma}}$. The temperature independent term in the energy sets an energy scale. Thus, to understand the effect of change in the energy on the dynamics, we should scale the energy by its temperature independent value. We find from Eq.21 and Eq.22 that the slopes of both $\Delta E^{ma}/A$ and $1/(\Delta E^{ma}*A^{\prime})$ are the same and equal to $B/A=B^{\prime}/A^{\prime}=\gamma T_{VFT}$. The values of $\gamma$ obtained from the temperature dependence of the macroscopic SOP (right panels in Fig.2) are given in Table 1. We observe that $\gamma T_{VFT}$ varies linearly with the fragility parameter $K_{VFT}$ across all three sets of systems, as demonstrated in Fig.4. Note that at the macroscopic level $\Delta E^{ma}$ as a function of SOP does not actually show a linear dependence (left panels of Fig.1) and Eq.21 is an approximate expression. However, as shown in the right panel of Fig.1, $1/\Delta E^{ma}$ as a function of SOP does show a linear behaviour validating Eq.22. Thus, we obtain the values of $A^{\prime}$ and $B^{\prime}$ from right panels in Fig.1 (given in Table.2 in the Appendix III). In the inset of Fig.4 we also plot $B^{\prime}/A^{\prime}$ against $\gamma T_{VFT}$. We find that all the points fall on the y=x line, thus confirming our analysis. The physical reason for $\gamma T_{VFT}$ being a good fragility metric needs further investigation. Earlier studiesZaccone_PNAS ; Zaccone_PRB have shown that the fragility of the modified LJ systems are linearly proportional to the product of thermal expansion coefficient and the glass transition temperature. It will be interesting to understand the correlation between the thermal expansion coefficient and the $\gamma$ parameter.
@@
Figure 4: The main panel displays $\gamma T_{\mathrm{VFT}}$ plotted against $K_{\mathrm{VFT}}$ for three sets of systems: LJ, WCA, and modified LJ $(q,p)$ system. Each set exhibits a clear linear relationship, indicating a consistent correlation between the parameters. Inset: The scaled slope of the macroscopic energy barrier plot, $B^{\prime}/A^{\prime}$, derived from the $1/\Delta E^{\mathrm{ma}}$ versus $1/\Delta((\beta\Phi)^{-1})$ plot in the right panel of Fig.1, is plotted against $\gamma T_{\mathrm{VFT}}$. All data points lie on the $y=x$ line, demonstrating that the two quantities are numerically identical, validating our analysis.
@@
Table 1: Fitted values of the VFT temperature, $T_{\mathrm{VFT}}$, and fragility, $K_{\mathrm{VFT}}$, for various systems, obtained using Eq.11. The parameters $\gamma$ and $\gamma^{{}^{\prime}}$ represent the temperature dependence of our SOP (Eq.LABEL:gamma) at macroscopic and microscopic levels respectively, (right panel of Fig.2). Note that $(12,6)$ system is same as LJ at $\rho=1.2$
@@
Figure 5: Left panels: Microscopic energy barrier, $\Delta E$, plotted as a function of inverse microscopic structural order parameter, $1/\Delta((\beta\phi)^{-1})$, for three sets of systems: (a)LJ and (c)WCA, each examined at four different densities, and (e) modified LJ $(q,p)$ system. Right panels: Scaled microscopic energy barrier, $\Delta E/A$, plotted against $1/\Delta((\beta\phi)^{-1})$ for the same systems shown in the left panels. The colour coding in the right panels matches that of the left panels.
@@
Next we study the validity of the same analysis for the microscopic $\Delta E$. In Fig.2(left panels) we plot the $\Delta E$ vs $1/\Delta((\beta\phi)^{-1})$. Notably, unlike at the macroscopic level where the energy barrier increases more than linearly with the structural order parameter(SOP) (left panels of Fig.1), at the microscopic level, $\Delta E$ as a function of SOP (left panels of Fig.2) exhibits a linear behavior. A similar linear trend is also seen in the $\Delta E$ vs. modified SOP plot (left panels of Fig.5). In the calculation of the microscopic energy barrier, it is assumed that the dynamics is given by $P_{R}^{-1}$ (see Appendix II). The fact that $\Delta E^{ma}$ grows more strongly compared to $\Delta E$ can be attributed to the fact that $P_{R}^{-1}$ is related to a shorter time dynamics which does not have the information of the cooperativity in the dynamics thus grows slower than $\tau_{\alpha}$, an observation previously reported by Tah et alTah2022_jcp . However, as discussed before, the approximate behaviour at the macroscopic and microscopic levels remains similar. For the LJ and the WCA system, we find that the rate of growth of the energy barriers follow the fragility, whereas for the modified LJ $(q,p)$ system, the behaviour is the reverse. We next plot in Fig.5 (right panels) the scaled energy $\Delta E/A$, where the A is obtained by fitting the SOP dependence of the energy barrier to a linear form $\Delta E=A+B.\frac{1}{\Delta((\beta\phi)^{-1})}$. According to our previous analysis, $B/A(micro)$, the slope of the scaled microscopic energy barrier should be equal to $\gamma T_{VFT}$. However, as shown in Fig.6(a) that does not seem to be the case. In the inset of Fig.6(a) we also plot $B/A(micro)$ against $K_{VFT}$, and again no consistent trend is observed. This behaviour is in stark contrast to that of the macroscopic counterpart. The slope of the microscopic energy barrier varies significantly across systems and does not follow a universal trend. For instance, in the modified LJ $(q,p)$ system $B/A(micro)$ appears to track both $\gamma T_{VFT}$ and $K_{VFT}$. In contrast, the WCA systems, which exhibit the largest variation in fragility, show almost no change in the $B/A(micro)$. Interestingly, for the LJ systems, the slope of the scaled microscopic energy barrier varies inversely with both $\gamma T_{VFT}$ and $K_{VFT}$.
@@
Figure 6: (a)The main panel presents the scaled slope of the microscopic energy barrier plot, $B/A(micro)$, obtained from the slope of the $\Delta E/A$ versus $1/\Delta((\beta\phi)^{-1})$ plot (right panel of Fig.5), plotted against $\gamma^{{}^{\prime}}T_{\mathrm{VFT}}$ for three sets of systems: LJ, WCA, and modified LJ $(q,p)$ system. The data exhibit distinct behaviours across the systems. Inset: $B/A(micro)$ plotted against $K_{\mathrm{VFT}}$ shows a similar trend to the main plot, consistent with the linear relationship between $\gamma T_{\mathrm{VFT}}$ and $K_{\mathrm{VFT}}$ (demonstrated in Fig.4). (b) The peak value of the Spearman rank correlation between the coarse grained SOP and the dynamics, $C_{R}((\beta\phi)^{-1},\mu)$, obtained from Fig.7, plotted against $B/A(micro)$.
@@
Thus, although the macroscopic and microscopic energy barriers appear to exhibit qualitatively similar variation with the SOP, a more detailed analysis reveals key differences. The scaled slope of the energy barrier as a function of SOP quantifies how sensitively the barrier and therefore, the dynamics responds to structural changes. A higher slope indicates that even small variations in SOP lead to significant changes in the energy barrier and dynamics, reflecting a stronger coupling between structure and dynamics.
@@
Accordingly, the macroscopic and microscopic energy barriers imply different strengths of structure-dynamics correlation. This discrepancy may originate from the nature of the analyses: macroscopic approaches average over many particles, thereby smoothing out local fluctuations and heterogeneities. In contrast, microscopic analysis preserves particle-level detail and captures spatially localised structural variations and dynamic heterogeneity. As a result, macroscopic models may overlook subtle but crucial features that govern dynamic behaviour.
@@ IV.2 Structure-Dynamics correlation via isoconfigurational study
As discussed before, in order to apply an independent method to study structure-dynamics correlation, we perform the study using isoconfiguration runs. This is a powerful technique which was proposed by Harrowell and co-workers Cooper2008_nature to extract the effect of any structural order parameter on the dynamics. To enable a meaningful comparison across systems with different densities, we select temperatures such that the $\alpha$-relaxation time is approximately the same for each system. The working temperatures used for comparing different densities are listed in Table 3 in Appendix IV. We then compute the Spearman rank correlation between the SOP at the initial time and the particle mobility measured at different times(Fig.7). As observed earlier by us Sharma2023_jcp and also discussed in detail in the literature Tong2018_prx ; Tong2019_Nature ; Tong2020_prl ; Mei2022_localvolume ; yethiraj_pnas2025 ; Tanaka2025_review , the correlation between an order parameter and the dynamics when the order parameter is calculated at a single particle level is high at short times and then drops at longer times. However, the correlation with coarse grained order parameter increases at long times. The coarse graining of a parameter $X$ for the $i^{th}$ particle, when coarse grained over a length scale $L$, is defined as $\bar{X}_{i}(L)=\sum_{j}X_{j}P(r_{j}-r_{i})/\sum_{j}P(r_{j}-r_{i})$, where $P(x)$ = $exp(-x/L)$, under the hypothesis that the influence of the neighbourhood decays exponentially with distance. Studies have shown that there is an optimal value of the coarse graining length where the correlation is maximum Tong2018_prx ; Tong2019_Nature ; Tong2020_prl ; Tanaka2025_review ; Sharma2023_jcp . The SOP is coarse-grained over this length, $L=4$ for LJ, and modified LJ $(q,p)$ system and $L=3$ for WCA systems in such a way that it yields maximum correlation. Note that, as shown in the left panels of Fig.15 in Appendix IV, although the value of the correlation changes with $L$, the nature of the plot remains similar.
Figure 7: Spearman rank correlation,$C_{R}((\beta\phi)^{-1},\mu)$, between the coarse-grained structural order parameter, $(\beta\phi)^{-1}$, and particle mobility, $\mu$, is plotted as a function of time scaled by the $\alpha$-relaxation time ($t/\tau_{\alpha}$) for three sets of systems: (a) LJ system at four densities: with coarse-graining length $L=4$; (b) WCA system at the four densities with $L=3$; (c) modified LJ $(q,p)$ system with $L=4$.
@@
Our analysis reveals that the structure-dynamics correlation is not always related to fragility. However, we find a strong relationship between the structure–dynamics correlation and the slope of the scaled microscopic energy barrier across all three sets of systems (Fig.6(a)): i) For the LJ system, both the slope and the structure–dynamics correlation decrease with increasing fragility. ii) For the WCA system, despite exhibiting the largest variation in fragility, both the slope and the structure–dynamics correlation remain nearly constant and appear independent of fragility. iii) For the modified LJ $(q,p)$ system, both the slope and the structure–dynamics correlation increase with fragility. Fig.6(b) shows the peak value of the structure-dynamics correlation $C_{R}$ against the scaled microscopic energy barrier $B/A(micro)$, which quantitatively underscores this observation.
@@
These findings allow us to conclude that the slope of the scaled microscopic energy barrier is closely linked to the structure-dynamics correlation. In contrast, the slope of the scaled macroscopic energy barrier, given by $\gamma T_{VFT}$ does follow fragility but not the structure-dynamics correlation. This implies that the structure–dynamics correlation is not trivially connected to fragility. However, phenomenologically, fragility has been linked to dynamic heterogeneity, with more fragile systems typically exhibiting stronger heterogeneity in dynamics Sastry2001_nature ; Ediger2000_Physchem . This raises an important and open question: What fundamental mechanism governs the system dependent trends in structure-dynamics correlation, and how is this related to dynamic heterogeneity?

### v4

@@ IV.1 The temperature and SOP dependence of the free energy barrier and its correlation with fragility
From Eq.18 we get that $A^{\prime}=\frac{K_{VFT}}{T_{VFT}}$ and $B^{\prime}=\gamma.K_{VFT}$, which shows that the absolute slope $B^{\prime}$ of the inverse of the free energy barrier is proportional to fragility. From Eq.19 we get A = $\frac{T_{VFT}}{K_{VFT}}$ and B = $\gamma.\frac{T_{VFT}^{2}}{K_{VFT}}$ and the slope just like that for temperature (see Appendix III) is again not proportional to but inversely proportional to the fragility. For both the LJ and WCA systems, fragility increases with density, accompanied by a corresponding rise in $T_{VFT}$ as the systems at higher densities operate in higher temperature regimes. This trend explains why with increasing fragility, we observe in Fig.1(left panels) and Fig.2 (left panels) a sharper increase in the free energy barrier, which is approximated by a linear expression given by Eq.19 with slope $\gamma.\frac{T_{VFT}^{2}}{K_{VFT}}$. For the modified LJ $(q,p)$ system, the fragility is higher for the $(8,5)$ system, which has the lowest value of $T_{VFT}$, and this is the reason the growth of the free energy barrier for this system is low. Thus, in cases where the growth of the free energy barrier follows fragility, it is because for more fragile systems the $T_{VFT}$ is higher.
@@
Note that AA($A^{\prime}$) describes the part of the activation free energy(inverse energy) which is independent of the temperature and also the structural order and BB($B^{\prime}$) describes the sensitivity of the dynamics to change in the structural order. To do a meaningful comparison of the free energy barrier and also the dynamics of different systems, we need to scale the free energy barrier by the temperature independent value. Thus we find that instead of BB($B^{\prime}$) it is the scaled quantity B/AB/A($B^{\prime}/A^{\prime}$) that describes the sensitivity of the dynamics to change in structural order. Interestingly, we find that $B/A=B^{\prime}/A^{\prime}=\gamma T_{VFT}$. The values of $\gamma$ obtained from the temperature dependence of the macroscopic SOP (right panels in Fig.2) are given in Table 1. We observe that $\gamma T_{VFT}$ varies linearly with the fragility parameter $K_{VFT}$ across all three sets of systems, as demonstrated in Fig.4. Note that at the macroscopic level $\Delta E^{ma}$ as a function of inverse SOP does not actually show a linear dependence (left panels of Fig.1) and Eq.19 is an approximate expression. However, as shown in the right panels of Fig.1, $1/\Delta E^{ma}$ as a function of SOP does show a linear behaviour, validating Eq.18. Thus, we obtain the values of $A^{\prime}$ and $B^{\prime}$ from the right panels in Fig.1 (given in Table.2 in Appendix III). In the inset of Fig.4 we also plot $B^{\prime}/A^{\prime}$ against $\gamma T_{VFT}$. We find that all the points fall on the y=x line, thus confirming our analysis. The physical reason for $\gamma T_{VFT}$ being a good fragility metric needs further investigation. Earlier studies Zaccone_PNAS ; Zaccone_PRB have shown that the fragility of the modified LJ systems is linearly proportional to the product of the thermal expansion coefficient and the glass transition temperature. It will be interesting to understand the correlation between the thermal expansion coefficient and the $\gamma$ parameter.
@@
Figure 4: The main panel displays $\gamma T_{\mathrm{VFT}}$ plotted against $K_{\mathrm{VFT}}$ for three sets of systems: LJ, WCA, and modified LJ $(q,p)$ system. Each set exhibits a clear linear relationship, indicating a consistent correlation between the parameters. Inset: The scaled slope of the macroscopic free energy barrier plot, $B^{\prime}/A^{\prime}$, derived from the $1/\Delta E^{\mathrm{ma}}$ versus $1/\Delta((\beta\Phi)^{-1})$ plot in the right panels of Fig.1, is plotted against $\gamma T_{\mathrm{VFT}}$. All data points lie on the $y=x$ line, demonstrating that the two quantities are numerically identical, validating our analysis.
@@
Table 1: Fitted values of the VFT temperature, $T_{\mathrm{VFT}}$, and fragility, $K_{\mathrm{VFT}}$, for various systems, obtained using Eq.11. The parameters $\gamma$ and $\gamma^{{}^{\prime}}$ represent the temperature dependence of our SOP (Eq.LABEL:gamma) at macroscopic and microscopic levels, respectively (right panels of Fig.2). Note that $(12,6)$ system is same as LJ at $\rho=1.2$
@@
Figure 5: Left panels: Microscopic free energy barrier, $\Delta E$, plotted as a function of inverse microscopic structural order parameter, $1/\Delta((\beta\phi)^{-1})$, for three sets of systems: (a)LJ and (c)WCA, each examined at four different densities, and (e) modified LJ $(q,p)$ system. Right panels: Scaled microscopic free energy barrier, $\Delta E/A$, plotted against $1/\Delta((\beta\phi)^{-1})$ for the same systems shown in the left panels. The colour coding in the right panels matches that of the left panels.
@@
Next we study the validity of the same analysis for the microscopic $\Delta E$. In Fig.2(left panels) we plot the $\Delta E$ vs $1/\Delta((\beta\phi)^{-1})$. Notably, unlike at the macroscopic level where the free energy barrier increases more than linearly with the structural order parameter(SOP) (left panels of Fig.1), at the microscopic level, $\Delta E$ as a function of SOP (left panels of Fig.2) exhibits a linear behaviour. A similar linear trend is also seen in the $\Delta E$ vs. modified SOP plot (left panels of Fig.5). In the calculation of the microscopic free energy barrier, it is assumed that the dynamics is given by $P_{R}^{-1}$ (see Appendix II). The fact that $\Delta E^{ma}$ grows more strongly compared to $\Delta E$ can be attributed to the fact that $P_{R}^{-1}$ is related to a shorter time dynamics which does not have the information of the cooperativity in the dynamics thus grows slower than $\tau_{\alpha}$, an observation previously reported by Tah et alTah2022_jcp . However, as discussed before, the approximate behaviour at the macroscopic and microscopic levels remains similar. For the LJ and the WCA system, we find that the rate of growth of the free energy barriers follows the fragility, whereas for the modified LJ $(q,p)$ system, the behaviour is the reverse. As discussed in the case of the macroscopic free energy barrier, to do a meaningful comparison of the free energy barrier of different systems, we need to scale it by its temperature independent value. We first fit the $\Delta E$ to a linear form $\Delta E=A+B.\frac{1}{\Delta((\beta\phi)^{-1})}$ (Fig.5 (left panels)). We find that the value of $A$ not only increases with the depth of the potential as in the case of the LJ $(q,p)$ system, but it also increases when the density of the system is increased, keeping the interaction potential the same. We next plot in Fig.5 (right panels) the scaled energy $\Delta E/A$. According to our previous analysis, $B/A(micro)$, the slope of the scaled microscopic free energy barrier should be equal to $\gamma^{\prime}T_{VFT}$. However, as shown in Fig.6(a), that does not seem to be the case. In the inset of Fig.6(a), we also plot $B/A(micro)$ against $K_{VFT}$, and again no consistent trend is observed. This behaviour is in stark contrast to that of the macroscopic counterpart. The slope of the microscopic free energy barrier varies significantly across systems and does not follow a universal trend. For instance, in the modified LJ $(q,p)$ system $B/A(micro)$ appears to track both $\gamma^{\prime}T_{VFT}$ and $K_{VFT}$. In contrast, the WCA systems, which exhibit the largest variation in fragility, show almost no change in the $B/A(micro)$. Interestingly, for the LJ systems, the slope of the scaled microscopic free energy barrier varies inversely with both $\gamma T_{VFT}$ and $K_{VFT}$.
@@
Figure 6: (a)The main panel presents the scaled slope of the microscopic free energy barrier plot, $B/A(micro)$, obtained from the slope of the $\Delta E/A$ versus $1/\Delta((\beta\phi)^{-1})$ plot (right panels of Fig.5), plotted against $\gamma^{{}^{\prime}}T_{\mathrm{VFT}}$ for three sets of systems: LJ, WCA, and modified LJ $(q,p)$ system. The data exhibit distinct behaviours across the systems. Inset: $B/A(micro)$ plotted against $K_{\mathrm{VFT}}$ shows a similar trend to the main plot, consistent with the linear relationship between $\gamma T_{\mathrm{VFT}}$ and $K_{\mathrm{VFT}}$ (demonstrated in Fig.4). (b) The peak value of the Spearman rank correlation between the coarse grained SOP and the dynamics, $C_{R}((\beta\phi)^{-1},\mu)$, obtained from Fig.7, plotted against $B/A(micro)$.
@@
Thus, although the macroscopic and microscopic free energy barriers appear to exhibit qualitatively similar variation with the SOP, a more detailed analysis reveals key differences. The scaled slope of the free energy barrier as a function of SOP quantifies how sensitively the barrier and therefore, the dynamics responds to structural changes. A higher slope indicates that even small variations in SOP lead to significant changes in the free energy barrier and dynamics, reflecting a stronger coupling between structure and dynamics.
@@
Accordingly, the macroscopic and microscopic free energy barriers imply different strengths of structure-dynamics correlation. This discrepancy may originate from the nature of the analyses: macroscopic approaches average over many particles, thereby smoothing out local fluctuations and heterogeneities. In contrast, microscopic analysis preserves particle-level detail and captures spatially localised structural variations and dynamic heterogeneity. As a result, macroscopic models may overlook subtle but crucial features that govern dynamic behaviour.
@@ IV.2 Structure-Dynamics correlation via isoconfigurational study
As discussed before, in order to apply an independent method to study structure-dynamics correlation, we perform the study using isoconfiguration runs. This is a powerful technique which was proposed by Harrowell and co-workers Cooper2008_nature to extract the effect of any structural order parameter on the dynamics. To enable a meaningful comparison across systems with different densities, we select temperatures such that the $\alpha$-relaxation time is approximately the same for each system. The working temperatures used for comparing different densities are listed in Table 3 in Appendix IV. We then compute the Spearman rank correlation between the SOP at the initial time and the particle mobility measured at different times(Fig.7). As observed earlier by us Sharma2023_jcp and also discussed in detail in the literature Tong2018_prx ; Tong2019_Nature ; Tong2020_prl ; Mei2022_localvolume ; yethiraj_pnas2025 ; Tanaka2025_review , the correlation between an order parameter and the dynamics when the order parameter is calculated at a single particle level is high at short times and then drops at longer times. However, the correlation with coarse grained order parameter increases at long times. The coarse graining of a parameter $X$ for the $i^{th}$ particle, when coarse grained over a length scale $L$, is defined as $\bar{X}_{i}(L)=\sum_{j}X_{j}P(r_{j}-r_{i})/\sum_{j}P(r_{j}-r_{i})$, where $P(x)$ = exp(x/L)exp(-x/L), under the hypothesis that the influence of the neighbourhood decays exponentially with distance. Studies have shown that there is an optimal value of the coarse graining length where the correlation is maximum Tong2018_prx ; Tong2019_Nature ; Tong2020_prl ; Tanaka2025_review ; Sharma2023_jcp ; sanket_lengthscale . The SOP is coarse-grained over this length, $L=4$ for LJ, and modified LJ $(q,p)$ system and $L=3$ for WCA systems in such a way that it yields maximum correlation. Note that, as shown in the left panels of Fig.15 in Appendix IV, although the value of the correlation changes with $L$, the nature of the plot remains similar.
[…]
@@
Our analysis reveals that the structure-dynamics correlation is not always related to fragility. However, we find a strong relationship between the structure–dynamics correlation and the slope of the scaled microscopic free energy barrier across all three sets of systems when we compare the plots in Fig.7 and the right panels of Fig.5: i) For the LJ system, both the slope and the structure–dynamics correlation decrease with increasing fragility. ii) For the WCA system, despite exhibiting the largest variation in fragility, both the slope and the structure–dynamics correlation remain nearly constant and appear independent of fragility. iii) For the modified LJ $(q,p)$ system, both the slope and the structure–dynamics correlation increase with fragility. Fig.6(b) shows the peak value of the structure-dynamics correlation $C_{R}$ against the scaled microscopic free energy barrier $B/A(micro)$, which quantitatively underscores this observation.
@@
These findings allow us to conclude that the slope of the scaled microscopic free energy barrier is closely linked to the structure-dynamics correlation. In contrast, the slope of the scaled macroscopic free energy barrier, given by $\gamma T_{VFT}$ does follow fragility but not the structure-dynamics correlation. This implies that the structure–dynamics correlation is not trivially connected to fragility. However, phenomenologically, fragility has been linked to dynamic heterogeneity, with more fragile systems typically exhibiting stronger heterogeneity in dynamics Sastry2001_nature ; Ediger2000_Physchem . This raises an important and open question: What fundamental mechanism governs the system dependent trends in structure-dynamics correlation, and how is this related to dynamic heterogeneity?